OpenStudy (anonymous):

Write the complex number in the form a + bi sqrt6(cos 315° + i sin 315°) PLEASE HELP

OpenStudy (anonymous):

options are:

OpenStudy (anonymous):

@jim_thompson5910 could you help with this last one :/ ? Im so sorry

jimthompson5910 (jim_thompson5910):

what is the cosine of 315 degrees

OpenStudy (anonymous):

sqrt2/2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what about the sine of 315 degrees

OpenStudy (anonymous):

-1/sqrt2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

or -sqrt(2)/2

OpenStudy (anonymous):

sorry about that youre right ./.... typo @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so this means that we have this so far \[\Large \sqrt{6}\left(\cos(315) + i\sin(315)\right)\] \[\Large \sqrt{6}\left(\frac{\sqrt{2}}{2} - i*\frac{\sqrt{2}}{2}\right)\]

OpenStudy (anonymous):

so like (2-2i) sqrt3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how did you get that

OpenStudy (anonymous):

I have no clue ... I messed up on the way Im thinking it was either A or C but Im leaning towards A

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{6}\left(\cos(315) + i\sin(315)\right)\] \[\Large \sqrt{6}\left(\frac{\sqrt{2}}{2} - i*\frac{\sqrt{2}}{2}\right)\] \[\Large \sqrt{6}*\left(\frac{\sqrt{2}}{2}\right) - i*\sqrt{6}*\left(\frac{\sqrt{2}}{2}\right)\] \[\Large \frac{\sqrt{6}*\sqrt{2}}{2} - i*\frac{\sqrt{6}*\sqrt{2}}{2}\] \[\Large \frac{\sqrt{6*2}}{2} - i*\frac{\sqrt{6*2}}{2}\] \[\Large \frac{\sqrt{12}}{2} - i*\frac{\sqrt{12}}{2}\] \[\Large \frac{\sqrt{4*3}}{2} - i*\frac{\sqrt{4*3}}{2}\] \[\Large \frac{\sqrt{4}*\sqrt{3}}{2} - i*\frac{\sqrt{4}*\sqrt{3}}{2}\] \[\Large \frac{2*\sqrt{3}}{2} - i*\frac{2*\sqrt{3}}{2}\] \[\Large \sqrt{3} - i*\sqrt{3}\] \[\Large \sqrt{3} - \sqrt{3}*i\]

OpenStudy (anonymous):

Thank you !!!

jimthompson5910 (jim_thompson5910):

np